OpenStudy (anonymous):

Find x_1 and x_2 that also satisfy the given initial conditions. x_1' = x_1+x_2 x_2' = 3x_1 - 4x_2 Where, x_1(0) = -1 x_2(0) = 2 I got the system of equations down to the general solution x = C1exp(-t) + C2exp(-2t) but I'm getting the incorrect values for c1 and c2 for x_1 and x_2.

OpenStudy (anonymous):

\[x = e^{-t}\left(\begin{matrix}-7 \\ -7\end{matrix}\right) + e^{-2t}\left(\begin{matrix}6 \\ 9\end{matrix}\right)\] This is how the answer looks like, my c_1 and c_2 values are completely off.